Q is a point charge in vacuum. The electric potential 15.0 mm from Q is 450 V greater than the potential 10 mm from Q. What is the charge on Q?
Answer #1

Electric potential is given by:

V = k*Q/R

Now given that electric potential 15.0 mm from Q is 450 V greater than the potential 10 mm from Q.

which means

V2 = V1 - 450

V2 = potential at R2 (15.0 mm)

V1 = potential at R1 (10.0 mm)

So,

V1 - V2 = 450

k*Q/R1 - k*Q/R2 = 450

k*Q*(1/R1 - 1/R2) = 450

Q = 450*R1*R2/(k*(R2 - R1))

Q = 450*15*10^-3*10*10^-3/(9*10^9*(15 - 10)*10^-3)

Q = 1.5*10^-9 C = 1.5 nC

Please Upvote.

Comment below if you have any query.